OpenStudy (anonymous):

Solve x2 + 10x + 7 = 0 by completing the square.

OpenStudy (anonymous):

Which equation is used in the process? (x + 10)2 = 93 (x + 5)2 = 32 (x + 5)2 = 18

OpenStudy (anonymous):

x^2+10x+7+18 -18=0 (x+5)^2-18=0 (x+5)^2=18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep anytime dont forget to fan me if you'd like :3